# SUBLLM: A Novel Efficient Architecture with Token Sequence Subsampling for LLM

### v3

@@ 1 Introduction
Recently in NLP field, the emergence of large language models (LLMs) marks a pivotal advancement in how machines understand and generate human language [4, 26, 48, 37, 40]. LLMs are pre-trained with huge amount of parameters on extensive data and gain extraordinary capabilities across a series of downstream tasks.
However, the training and inference of LLMs require huge resources including memory and computational power, challenging their widespread deployment [36, 50]. To alleviate this problem, previous studies have endeavored to compress LLMs through knowledge distillation [15, 16, 30], quantization [41, 11] and pruning [8, 22, 10]. Another stream of research accelerates the inference process through decoding optimization [5, 20, 33], continuous batching, and conditional computation [2, 9, 19]. Nevertheless, all above approaches focus on optimization in the post-pretraining phase and fall short of fundamentally mitigating the pre-training computational resource demands.
Standard Transformer architectures treat all tokens in a sequence equally. However, research [6, 14, 1] has shown that in natural language, some tokens play a more crucial role while others are less important. Inspired by this finding, if we can identify and selectively remove those less important tokens, we can significantly reduce computational demands. Moreover, focusing on the most crucial information by removing less important tokens could enhance training stability and accelerate convergence, potentially leading to improved modeling outcomes [28].
In this paper, we propose a novel and efficient architecture Subsampling-Upsampling-Bypass Large Language Model (SUBLLM), inheriting the structure of decoder-only LLM, which dynamically allocates computational resources for tokens according to their importance. Inspired by U-Net architecture [29], SUBLLM integrates subsampling and upsampling modules symmetrically between the Transformer blocks, reducing the computational cost while preserving the input sequence’s semantics. Specifically in the subsampling module, a scoring layer calculates each token’s importance as the criterion for token subsampling. Meanwhile, a balancer is adopted to adjust the distribution of the token-level score during training. Subsequently, the upsampling module recovers the subsampled sequences to their prior lengths for token prediction in language modeling. Moreover, to improve training stability and accelerate convergence speed, SUBLLM integrates a bypass module that performs a weighted sum of the upsampled token sequence and the original one. The experimental results compared with LLaMA [37] demonstrate the effectiveness of our proposed SUBLLM on model efficiency as well as performance maintenance. The main contributions of this work are summarized as follows:
@@ 2 Related Work
2.1 Training Accerlation
To reduce the computational cost in LLM training, a lot of work has been carried out from the perspective of reducing redundancy in text. Funnel-Transformer [6] uses strided mean pooling to gradually compress the sequence of hidden states in self-attention, and Fourier-Transformer [14] progressively subsamples hidden states with the Fast Fourier Transform operator. In contrast to these methods, our proposed SUBLLM subsamples the tokens to a shortened sequence and reconstructs them in the upsampling module, which not only effectively reduces computational burden, but also guarantees the performance.
Another line of work leverages conditional computation [2, 9, 1] to dynamically expend resources when needed. CoLT5 [1] uses conditional routing to decide whether a given token passes through a light branch or a heavy branch in feedforward and attention layers, so as to devote more resources to important tokens. Further, MoD [28] utilizes a static compute budget, using a per-block router to select tokens for computations, and optimizing FLOP usage by choosing between self-attention and MLP blocks or a residual connection. Compared with CoLT5, MoD focuses on the decoder-only setting and overcomes the non-causal nature of the top-k operation. Our method can also be seen as conditional computation, which dynamically allocates the computational resources to tokens tailored to their importance.
Mamba [13] introduces selective state space models, allowing the model to filter out irrelevant information based on the input, enhancing content-based reasoning. Some other work proposes linear attention [39, 17, 47] for training efficiency. These techniques, which do not use the original attention mechanism, are thus incompatible with existing decoder-only LLMs. Additionally, CEPE [45] devises parallel encoding with chunk encoders processing long context in parallel, and the cross-attention layers in the decoder leverage the additional encoder, which increases the throughput and decreases the memory cost. Compared to this, our proposed SUBLLM places subsampling and upsampling modules at symmetrical positions between Transformer blocks, and thus is compatible with current LLMs and reduces token sequence length.
@@ 2.2 Inference Acceleration
The parameters that define Large Language Models (LLMs) are both vast and complex, leading to an ever-increasing demand for computational power and memory capacity. To address these challenges, prior research has primarily concentrated on developing more lightweight models derived from their heavier pre-trained counterparts. This strategy aims to reduce the intensive memory requirements and enhance the speed of inference processes. Key techniques employed in this endeavor include knowledge distillation [15, 16, 30, 23, 12], quantization [41, 11, 21, 43, 46] and pruning [8, 22, 10, 35].
Beyond the sheer scale of Large Language Models (LLMs), a significant challenge impacting inference speed is the autoregressive decoding process. The language model decodes text sequentially, necessitating $K$ serial iterations to generate $K$ tokens. This step-by-step processing not only slows down the response time but also becomes a major bottleneck due to memory bandwidth limitations. Numerous efforts have been made to optimize the decoding process. Speculative decoding [20, 5, 33] samples from multiple tokens generate from efficient approximation model concurrently as speculative prefixes for the large model. Other methods optimize the decoding process with reference [42, 24]. For example, LLMA [42] selects text spans from the reference document and checks their appropriateness in parallel within a single decoding step. Recently, CoDA [19] integrates sparse activation and a set of new parameters into an existing densely pre-trained model, along with a brief training phase for inference efficiency improvement. In this work, our proposed new architecture SUBLLM is not mutually exclusive with the inference acceleration method above. On the contrary, SUBLLM can also leverage the previously mentioned strategies to expedite the inference process and reduce memory cost.
@@ 3 SUBLLM
As illustrated in Figure 1, the proposed SUBLLM model is based on the decoder-only LLM architecture. To manage the number of tokens processed, subsampling and upsampling modules are integrated into the Transformer blocks. The operation proceeds as follows: Initially, the model uses several Transformer blocks to process the full sequence, capturing a comprehensive token sequence representation. Subsampling modules are then introduced, which sequentially eliminate the less critical tokens, thereby reducing the sequence length required for processing. The highest level of sequence compression occurs in the network’s middle blocks. Subsequent to this, upsampling modules are employed to incrementally reinstate the sequence length. These modules merge the shorter, processed sequences with the original sequences before subsampling, restoring them to their full lengths. This mechanism allows the decoder-only model to operate as a language model, generating tokens sequentially, which is characteristic of language models where the input and output sequence lengths are identical. Additionally, we have incorporated bypass connection modules after the upsampling process to utilize each pre-subsampling embedding, assisting to improve the learning process from subsampling to upsampling. Our subsequent experiments confirm that this approach significantly improves convergence efficiency.
@@ 3.1 Learnable Subsampling Module
Token Selection The indexes of the kept tokens are determined by their importance values. To evaluate the importance of each token, a scoring layer predicts the token-level scalar importance value
$\textbf{w}=w_{1},\dots,w_{N}$, also known as weight:
@@ 4.1 Settings
We use SlimPajama [31] as the pre-training corpus, which includes CommonCrawl, C4, Wikipedia, GitHub, StackExchange, ArXiv, and Book. We sample each dataset according to the original SlimPajama data source proportions.
@@
In this study, we adopt LLaMA as the baseline model and train a version with 1.3B parameters, as presented in the main results. This version shares the same architecture and tokenizer as LLaMA2 [38], with a context window size of 4096. The training involves processing a total of 138B tokens. We conduct the training using the Fairseq [25] framework with bf16 precision, employing Flash Attention [7] to expedite the process. For SUBLLM, the training configuration is identical, except for the addition of 8192 extra parameters, which encompass the parameters of two scoring layers and two bypass modules. Compared to the original parameters, the additional parameters introduced by SUBLLM are negligible. SUBLLM undergoes two subsampling stages, each with a retention ratio of approximately 63.24% and a minimum retention ratio of 40% in the middle of the blocks. Additionally, comparisons for the 0.25B models under context windows of 2k to 8k are also discussed and analyzed in subsequent sections. This facilitates a thorough evaluation of SUBLLM’s performance across varying scales and conditions. Table 1 illustrates the structures of the 1.3B and 0.25B models. The 1.3B model, corresponding to the 24-block configuration, and the 0.25B model, corresponding to the 15-block configuration, both utilize two subsampling modules. The 1.3B model is trained on four nodes of 8 A100 GPUs, each with 80GB memory, while the 0.25B model is trained on one node. ZeRO [27] is utilized, configured with the ScaledAdam [44] optimizer, and Eden [44] learning rate scheduler is applied. At the beginning of training, we set $c_{min}$ and $c_{max}$ in the bypass module to 0.9 and 1.0 respectively. Then, we gradually decrease $c_{min}$ to 0.2 after 20k steps. The inference is performed on one GPU of A100.
@@ 4.2 Main Results
In evaluating the model performance of SUBLLM, we consider both the valid loss during pre-training and its performance on few-shot learning tasks. As shown in Table 2, SUBLLM’s valid loss is on par with that of LLaMA, indicating that its token prediction capabilities are comparable. For few-shot learning, we evaluate SUBLLM on 6 text classification datasets including sentiment classification (SST2 [32], Amazon and Yelp [49]), topic classification (DBpedia [18], AGNews [49]) and hate speech detection (Hate [3]). Despite some fluctuations in scores across different datasets, the overall performance of SUBLLM is broadly equivalent to LLaMA. Both findings above indicate the validity of the optimized architecture on the model performance in token prediction as well as few-shot in-context learning.
@@ 4.3 Ablation Study
We perform an ablation study on the 0.25B SUBLLM model to examine the effects of the bypass module on validation loss. The findings summarized in Table 3 show that SUBLLM with all enhancements achieves the lowest valid loss of 3.66. Changing the bypass module’s operation from a weighted sum to a standard residual connection increases the validation loss, which has been higher than LLaMA. It indicates the importance of the weighted integration. Completely removing the bypass module leads to a further increase in validation loss, which confirms the bypass module’s role in maintaining low valid loss by using intermediate token information. Overall, the bypass module significantly enhances learning efficiency.
@@ 6 Conclusion
In this study, we propose SUBLLM, a novel network architecture that utilizes text sequence redundancy and token significance to enhance training and decoding speeds while preserving few-shot learning capabilities. SUBLLM features an innovative subsampling mechanism allowing for customizable token retention ratios and includes a bypass module that significantly speeds up model convergence. Our findings indicate that the ScaledAdam optimizer supports this architecture by enhancing its convergence performance. This architecture is compatible with existing optimization methods within the LLaMA model family, ensuring wide applicability. Future research will investigate the impact of tokenizer compression ratio on SUBLLM to further understand token sequence subsampling.

### v5

@@ 1 Introduction
Recently in NLP field, the emergence of large language models (LLMs) marks a pivotal advancement in how machines understand and generate human language [5, 26, 37]. Pretrained with huge amounts of parameters on extensive data, LLMs gain extraordinary capabilities across a series of downstream tasks.
Though exhibiting remarkable potential in handling complex tasks, LLMs encounter challenges during training and inference. Firstly, the training process is extremely time-consuming, necessitating the processing of vast amounts of data. Secondly, they need a large amount of GPU memory and computational resources. These factors pose a challenge to their widespread deployment [36, 49].
To address these issues, several approaches have been proposed to accelerate inference and reduce computational costs. Techniques such as distillation [14, 21], quantization [10, 42], and pruning [9, 29]are employed, as well as decoding optimization [19, 32] and conditional computation [2, 18]. Additionally, many efforts are dedicated to training acceleration which often leads to inference acceleration too. Some focus on reducing text redundancy [1, 7, 12, 28] while others tackle the quadratic computational complexity of the Transformer’s self-attention mechanism by improving the attention mechanism [40, 45] and proposing new architectures [4, 11, 20, 25].
Drawing from pertinent research [1], natural language tokens in a sequence vary in importance. Selectively identifying and removing less significant tokens can significantly reduce computational demands. Moreover, this targeted approach to prioritizing key information has the potential to enhance training stability, accelerate convergence, and improve overall modeling performance [28].
In this paper, we propose a novel and efficient architecture Subsampling-Upsampling-Bypass Large Language Model (SUBLLM), inheriting the structure of decoder-only LLM, which dynamically allocates computational resources for tokens according to their importance. SUBLLM integrates subsampling and upsampling modules symmetrically between the Transformer blocks, reducing the computational cost while preserving the input sequence’s semantics. Specifically, in the subsampling module, a scoring layer calculates each token’s importance as the criterion for token subsampling. Meanwhile, a balancer is adopted to adjust the distribution of the token-level score during training. Subsequently, the upsampling module recovers the subsampled sequences to their prior lengths for token prediction in language modeling. Moreover, to improve training stability and accelerate convergence speed, SUBLLM integrates a bypass module that performs a weighted sum of the upsampled token sequence and the original one. The experimental results compared with LLaMA [37] demonstrate the effectiveness of our proposed SUBLLM on model efficiency as well as performance maintenance. The main contributions of this work are summarized as follows:
@@ 2 Related Work
2.1 Training Acceleration
To reduce the computational cost in LLM training, a lot of work has been carried out from the perspective of reducing redundancy in text. Funnel-Transformer [7] uses strided mean pooling to gradually compress the sequence of hidden states in self-attention, and Fourier-Transformer [12] progressively subsamples hidden states with the Fast Fourier Transform operator. Same to these methods, our proposed SUBLLM subsamples the tokens to a shortened sequence. Unlike these methods that reconstruct a sequence by repeating the reduced sequence and adding it back to the original for the final result, SUBLLM’s upsampling module takes a different approach. It interpolates between the original and subsampled sequences using token scores as weights, offering a more refined handling of sequence information.
Some other work leverages conditional computation to dynamically expend resources when needed. CoLT5 [1] uses conditional routing to decide whether a given token passes through a light branch or a heavy branch in feedforward and attention layers, so as to devote more resources to important tokens. Further, MoD [28] utilizes a static compute budget, using a per-block router to select tokens for computations, and optimizing FLOP usage by choosing between self-attention and MLP blocks or a residual connection. Our method can also be seen as conditional computation, which dynamically allocates the computational resources to tokens tailored to their importance.
Another type of work focuses on solving the inefficiency problem caused by the attention mechanism when transformers process sequences. Some works focus on improving the attention mechanism to increase the training efficiency [40, 45]. More recent efforts have introduced novel model architectures to overcome this limitation. RWKV [25] addresses limitations in Transformers by replacing quadratic QK attention with a linear scalar formulation. RecurrentGemma [4] combines linear recurrences with local attention to achieve high performance on language tasks with reduced memory usage and faster inference on long sequences. Mamba [11] introduces selective state space models, allowing the model to filter out irrelevant information based on the input, enhancing content-based reasoning. MEGALODON [20] introduces an enhanced MEGA architecture with several novel technical components designed to improve capability, efficiency, and scalability. While MEGALODON accelerates training for long sequence inputs with a 32K window length, its training speed at a 4K window length is slower than that of LLaMA. In contrast, our proposed SUBLLM builds upon the LLaMA model structure by incorporating subsampling modules to reduce sequence length. As a result, SUBLLM surpasses LLaMA in training efficiency, starting from a window length of 2K.
@@ 2.2 Inference Acceleration
The parameters that define LLMs are both vast and complex, leading to an ever-increasing demand for computational power and memory capacity. To address these challenges, prior research has primarily concentrated on developing more lightweight models derived from their heavier pre-trained counterparts. Key techniques employed in this endeavor include knowledge distillation [14, 21], quantization [10, 42] and pruning [9, 29]. However, these methods usually seek the balance between effect and inference acceleration at the cost of sacrificing model performance. While the novel model structure SUBLLM we proposed can not only accelerate the inference process but also maintain competitive performance.
Beyond the sheer scale of LLMs, a significant challenge impacting inference speed is the autoregressive decoding process. The language model decodes text sequentially, requiring K serial iterations to generate K tokens. This step-by-step processing not only delays the response time but also turns into a major bottleneck because of the limitations of memory bandwidth. Numerous efforts have been made to optimize the decoding process. For instance, speculative decoding [19, 22, 32] samples from multiple tokens generate from efficient approximation model concurrently as speculative prefixes for the large model. LLMA [43] accelerates the decoding process by identifying and utilizing overlapping text spans between the LLM output and the reference document. Medusa [6] expands the model’s predictive capabilities through additional heads and a specific tree-based attention mechanism. MInference [15] accelerates the pre-filling stage by leveraging dynamic sparse attention with spatial aggregation patterns. YOCO [34] optimizes inference efficiency by reusing global KV caches through cross-attention. Reducing the KV cache is also an effective method for accelerating inference [17, 23, 47].
In this work, our proposed new architecture SUBLLM is not mutually exclusive with the inference acceleration method above. On the contrary, SUBLLM can also leverage the previously mentioned strategies to expedite the inference process and reduce memory cost.
@@ 3 SUBLLM
As illustrated in Figure 1, the proposed SUBLLM model is based on the decoder-only LLM architecture. To manage the number of tokens processed, subsampling and upsampling modules are integrated into the Transformer blocks. The operation proceeds as follows: Initially, the model uses several Transformer blocks to process the full sequence, capturing a comprehensive token sequence representation. Subsampling modules are then introduced, which sequentially eliminate the less critical tokens, thereby reducing the sequence length required for processing. The highest level of sequence compression occurs in the network’s middle blocks. Subsequent to this, upsampling modules are employed to reinstate the sequence length. These modules merge the shorter, processed sequences with the original sequences before subsampling, restoring them to their full lengths. This mechanism allows the decoder-only model to operate as a language model, generating tokens sequentially, which is characteristic of language models where the input and output sequence lengths are identical. Additionally, we have incorporated bypass connection modules after the upsampling process to utilize each pre-subsampling embedding, assisting to improve the learning process from subsampling to upsampling. Our subsequent experiments confirm that this approach significantly improves convergence efficiency.
@@ 3.1 Learnable Subsampling Module
Token Selection The indexes of the kept tokens are determined by their importance values. To evaluate the importance of each token, a scoring layer predicts the token-level scalar importance value $\textbf{w}=w_{1},\dots,w_{N}$, also known as weight:
@@ 4.1 Settings
We use SlimPajama [30] as the pre-training corpus, which includes CommonCrawl, C4, Wikipedia, GitHub, StackExchange, ArXiv, and Book datasets, sampled according to SlimPajama’s original proportions.
@@
We adopt LLaMA2 [38] as the baseline, training 1.3B [41] and 0.25B parameter versions. SUBLLM shares the same training configuration but introduces only 8,192 additional parameters for the 1.3B model and 4,096 for the 0.25B model. Each model is trained with 100 times the number of its parameters in tokens and has versions with context window sizes of 2K, 4K, and 8K. Eden [44] is used for the learning rate schedule, ScaledAdam [44] as the optimizer, and ZeRO [27] to enhance training efficiency and optimize resource utilization. Training is conducted with bf16 precision using Fairseq [24], with Flash Attention [8] to accelerate the process. More details are in the supplementary material [39].
@@ 4.2 Main Results
In evaluating the model performance of SUBLLM, we consider both the valid loss during pre-training and its performance on few-shot learning tasks. As shown in Table 2, SUBLLM’s valid loss is on par with that of LLaMA, indicating that its token prediction capabilities are comparable. For few-shot learning, we evaluate SUBLLM on 6 text classification datasets including sentiment classification (SST2 [31], Amazon and Yelp [46]), topic classification (DBpedia [16], AGNews [46]) and hate speech detection (Hate [3]). Despite some fluctuations in scores across different datasets, the overall performance of SUBLLM is broadly equivalent to LLaMA. Both findings above indicate the validity of the optimized architecture on the model performance in token prediction as well as few-shot in-context learning. The results of the model on other benchmarks [13, 35, 48] can be found in Table 8 in the supplementary materials [39].
@@ 4.3 Ablation Study
We perform an ablation study on the 0.25B SUBLLM model to examine the effects of the bypass module on validation loss. The findings summarized in Table 3 show that SUBLLM with all enhancements achieves the lowest valid loss of 3.66. Changing the bypass module’s operation from a weighted sum to a standard residual connection increases the validation loss, which is higher than LLaMA. This results demonstrates the importance of weighted integration. Completely removing the bypass module leads to a further increase in validation loss, which confirms the bypass module’s role in maintaining low valid loss by using intermediate token information. Overall, the bypass module significantly enhances learning efficiency.
@@ 6 Conclusion
In this study, we propose SUBLLM, a novel network architecture that utilizes text sequence redundancy and token significance to enhance training and decoding speeds while preserving few-shot learning capabilities. SUBLLM features an innovative subsampling mechanism allowing for customizable token retention ratios and includes a bypass module that significantly speeds up model convergence. Our findings indicate that the ScaledAdam optimizer supports this architecture by enhancing its convergence performance. This architecture is compatible with existing optimization methods within the LLaMA model family, ensuring wide applicability. Future research will investigate the impact of sequence compression ratios on SUBLLM to further understand token sequence subsampling, as well as further validate the model’s scalability.
Appendix A Training Details
Table 1 illustrates the structures of the SUBLLM models used in the experiments. The trained 0.25B models corresponds to the second row, which has a 15-block configuration, while the 1.3B model corresponds to the third row, which has a 24-block configuration. Both models utilize two subsampling modules with each retention ratio of approximately 63.24%, dropping to a minimum of 40% midway through the blocks. Table 7 lists the configurations for the LLaMA and SUBLLM model architectures.
The 1.3B model is trained on four nodes of 8 A100 GPUs, each with 80GB memory, while the 0.25B model is trained on one node. We use ScaledAdam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$. At the beginning of training, we set $c_{min}$ and $c_{max}$ in the bypass module to 0.9 and 1.0 respectively. Then, we gradually decrease $c_{min}$ to 0.2 after 20k steps. The inference is performed on one GPU of A100.
Table 7: Model configurations of 1.3B and 0.25B models.
Configuration 0.25B 1.3B
num_hidden_layers 15 24
hidden_size 1024 2048
intermediate_size 4096 5504
num_attention_heads 16 16
tie_word_embeddings false false
rope_theta 10000.0 10000.0
Appendix B Module Details
Figure 4 illustrates the computation process of subsampling, upsampling, and bypass. To enhance readability and facilitate understanding, we have consolidated the operations from Equation 12 to Equation 15 within the subsampling module.
Figure 4: A demonstration of the computation of subsampling, upsampling and bypass modules. From bottom to top, the first dashed box represents the learnable subsampling module, the second dashed box represents the upsampling module, and the third dashed box represents the bypass module. Sampling denotes performing sampling with replacement from the weights of the tokens not selected, and these weights are used to be subtracted from the weights of the selected tokens.
Appendix C Model Performance
Figure 5: Valid loss comparison of 1.3B models under 4K and 8K Settings.
Performance on Benchmarks
We report results for 1.3B SUBLLM and LLaMA models with 4k context length across three benchmarks, MMLU (5-shot), BBH (3-shot) and AGIEval (5-shot). As shown in Table 8, the results of both models on the benchmarks are close to random guessing due to the insufficient amount of training data. However, SUBLLM performs slightly better than LLaMA.
Table 8: Results for 1.3B SUBLLM and LLaMA models with 4K context length across three benchmarks.
Model MMLU BBH AGIEval
LLaMA 26.23 23.70 16.76
SUBLLM 26.41 24.17 17.64
Results on Valid Loss
We report the valid loss of the 1.3B SUBLLM and LLaMA models trained with 4k and 8k context lengths. The SUBLLM is configured with twice subsampling and a total retention ratio of 40%. As shown in Figure 5, the valid loss of SUBLLM is significantly lower than that of LLaMA under different context window lengths. This indicated the potential of our proposed model architecture SUBLLM in accelerating training while maintaining model performance. The gap in valid loss between LLaMA and SUBLLM widens as the training context window length increases. This might imply the potential of SUBLLM in handling long sequences. We leave this exploration for future work.
Appendix D Limitation
Our work faced two primary limitations that affected the scalability of our proposed architecture. First, due to limited computational resources, we were only able to train models with a maximum parameter size of 1.3 billion. This restriction prevented us from exploring the full potential of larger models, which could offer improved performance on more complex tasks. Second, the amount of data we used for training was also constrained. In our experiments with the 0.25B model, we found that using data equivalent to one hundred times the model’s parameters was sufficient for convergence. We applied the same approach to the 1.3B models, which meant we could not leverage the entire Slimpajama dataset. Together, these limitations hindered our ability to fully assess the model’s scalability. In future work, we will focus on validating the scalability of the proposed model architecture with larger models and more extensive datasets.